OpenStudy (anonymous):

given that y = 2x^-2, find dy/dx from first principles

OpenStudy (tanner23456):

bring the exponent to the front and multiply by the 2, then subtract 1 from the exponent. dy/dx = -4x^-3

OpenStudy (anonymous):

no, from first princicples means the whole dy/dx=lim(h->0) [f(x+h)-f(x)]/h

OpenStudy (anonymous):

you did the short way

OpenStudy (tanner23456):

Oh, woops. I forgot about that.

OpenStudy (tanner23456):

I haven't used that since calc 1

OpenStudy (badhi):

\[\begin{align*} \frac{dy}{dx}&=\lim \limits_{\Delta x\to 0}\frac{f(x+\Delta x)-f(x)}{\Delta x}\\ &=\lim \limits_{\Delta x\to 0}\frac{2(x+\Delta x)^{-2}-2x^{-2}}{\Delta x}\\ &=2\lim \limits_{\Delta x\to 0}\frac{x^{2}-(x+\Delta x)^2}{x^2(\Delta x )(x+\Delta x)^2}\\ &=\frac{2}{x^2}\lim \limits_{\Delta x\to 0}\frac{-\Delta x(2x+\Delta x)}{(\Delta x )(x+\Delta x)^2}\\ &=\frac{2}{x^2}\lim \limits_{\Delta x\to 0}\frac{-(2x+\Delta x)}{(x+\Delta x)^2}\\ &=\frac{2}{x^2}\frac{-2x}{x^2}\\ &=-4x^{-3} \end{align*}\]